1) It takes 62.25 ml of .04 M Na2S2O3 solution to completely react with the iodine present in a 175 ml idoine solution. How many grams of iodine (I2) were initially present in the sample and what is the final concentration (in mol/L) of I- ions in the solution
Answer #1

Equation of the reaction: 2 Na2S2O3 + I2 Na2S4O6 + 2 NaI

Moles of Na2S2O3 = molarity * Volume = 0.04 * 62.25 * 10-3 = 2.49 * 10-3 moles

from the equation, moles of I2 = moles of Na2S2O3 / 2 = 2.49 / 2 * 10-3 moles = 1.245 * 10-3 mol

Mass of I2 = moles * molar mass = 1.245 * 10-3 * 254 = 0.316 g

From the equation, [Na2S2O3] = [NaI] = [I-] = 0.04 M